OpenStudy (anonymous):

If f(x) = |x − 5| + 2, find f(3). 10 6 4 0

OpenStudy (anonymous):

4

HanAkoSolo (jamierox4ev3r):

@larryboxaplenty …please provide a process, then your answer. Keep in mind that providing answers repeatedly without a process first can result in consequences, thank you

OpenStudy (anonymous):

(x-5) = (3-5) = -2 abs(-2) = 2 You always make numbers positive. 2+2 = 4

HanAkoSolo (jamierox4ev3r):

basically, what you would do is plug 3 into the x-variable of the equation. It would look like this: f(3)=|3-5|+2

HanAkoSolo (jamierox4ev3r):

the absolute value of 3-5 is 2, because absolute value is always positive. add 2, and your solution is 4.

HanAkoSolo (jamierox4ev3r):

@Nylyak let me know if you understand, and feel free to ask me any questions :)

OpenStudy (anonymous):

Thank you @Jamierox4ev3r and @larryboxaplenty for explaining it. <3